OpenStudy (anonymous):

what is the equation of a line that passes through (0,0) and (5,3)??

OpenStudy (anonymous):

y = x5/3

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

or "rise (change in y values) / run (change in x values)"

OpenStudy (anonymous):

isnt that the slope tho? i need the equation the answers in the multiple choice are: a) 3y-5x=0 b) 5x+3y=0 c) x+xy=8 d) 5y-3x=0 e) None of the above

OpenStudy (anonymous):

ok...you've got the answer now rearrange the equation

OpenStudy (anonymous):

come on try it .....it can't hurt

OpenStudy (anonymous):

but you have a division symbol in your and i dont use division or multiplication

OpenStudy (anonymous):

i mean the multiple choice not me

OpenStudy (anonymous):

rearrange it....multiply both sides by 3...then move the xterm to the other side ...what do you get

OpenStudy (anonymous):

i just dont understand your equation x5/3?

OpenStudy (anonymous):

is that a fraction or division?

OpenStudy (anonymous):

a fraction and division are the same thing 5/3 means 5 parts out of 3 and 5 divided by 3

OpenStudy (anonymous):

ok true would it be b) 5x+3y=0?

OpenStudy (abstracthaven):

d.

OpenStudy (anonymous):

so 5/3*3 = 3 and 3*y = 3y these are just sybolic representaions they essentially mean the same thing

OpenStudy (anonymous):

oh i meant d i got the axis mixed up

OpenStudy (anonymous):

sorry my original eq was wrong:y = x3/5

OpenStudy (anonymous):

and yes d is the correct answer

OpenStudy (anonymous):

ok perfect thanks for the help i have to do an assessment test for college wednesday and i forgot most of the stuff we learned :(

OpenStudy (abstracthaven):

its d. because 5y-3x=0 5y=3x y=3/5x

OpenStudy (anonymous):

don't worry...you'll sail through just remember start with the basic and you ca't go wrong